A manufacturer of cotton pins knows that 5 percent of his product is

defective. If he sells cotter pins in boxes of 100, and guarantee that not more than 4

pins will be defective, what is the approximate probability that a box will, fail to

meet the guaranteed quality.

b) Which distribution is used in (a)also give the reason .


1
Expert's answer
2021-06-02T07:25:43-0400

a) Using Bernulli formula:

"P(<=4) = 1 - P(0) - P(1) - P(2) - P(3) - P(4) = 1 - 0.95^{100} - 99\u00d70.05\u00d70.95^{99} - 97\u00d749\u00d70.05^2\u00d70.95^{98} - 95\u00d716\u00d797\u00d70.05^3\u00d70.95^{97} - 93\u00d794\u00d795\u00d74\u00d70.05^4\u00d70.95^{96} = 0.607."

b) It is discrete binomial distribution, because it represents the probability for no more than 4 fails in 100 trials, given a fail probability of 0.05 for each trial.